# Determination of the stably free cancellation property for orders

### v2

@@ 11.2. Primary decomposition and generators of unit groups of quotient rings
Let $\mathcal{O}_{C}$ be the (unique) maximal $\mathcal{O}$-order in $C$. This is equal to the integral closure of $\Gamma$ in $C$. By the going-up theorem each of the maximal ideals $\mathfrak{p}_{i}$ is of the form $\mathfrak{P}_{i}\cap\Gamma$ for some maximal ideal $\mathfrak{P}_{i}$ of $\mathcal{O}_{C}$. Let $\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\}$ be the set of maximal ideals of $\mathcal{O}_{C}$ with $\mathfrak{g}\mathcal{O}_{C}\subseteq\mathfrak{P}_{i}$. Then
@@
Let $\Lambda$ be an $\mathcal{O}$-order in a finite-dimensional semisimple $K$-algebra $A$ and let $\mathfrak{h}$ be a proper full two-sided ideal of $\Lambda$. Then there exists an algorithm that determines generators of $(\Lambda/\mathfrak{h})^{\times}$.
@@
Let $C$ denote the centre of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order contained in $C$. Let $\mathfrak{g}=\mathfrak{h}\cap C$. Then $\mathfrak{g}$ is an ideal of an $\mathcal{O}$-order $\Gamma$, which may be strictly contained in $\mathcal{O}_{C}$. In the case that $\mathfrak{h}$ is equal to $\mathfrak{f}$, a two-sided ideal of a maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$, the desired algorithm is described in [BHJ22, §6.6] (this is a minor modification of [BB06, §3.4–3.7]). This relies crucially on the computation of the primary decomposition of $\mathfrak{g}$, which can be accomplished using work of Bley–Endres [BE05] under the assumption that $\Gamma=\mathcal{O}_{C}$ (which must hold when $\mathfrak{h}=\mathfrak{f}$). In fact, the computation of the primary decomposition of $\mathfrak{g}$ is the only obstacle to generalising the algorithm to arbitrary choices of $\mathfrak{h}$. By removing the restriction that $\Gamma=\mathcal{O}_{C}$ on the computation of primary decompositions in Corollary 11.3, we thereby obtain an algorithm for the computation of generators of $(\Lambda/\mathfrak{h})^{\times}$ for arbitrary proper full two-sided ideals $\mathfrak{h}$ of $\Lambda$. ∎
@@ 12.1. Setup
Henceforth assume that $A$ admits a decomposition $A=A_{1}\oplus A_{2}$ of $K$-algebras such that $A_{1}$ satisfies the Eichler condition. In applications, we can always assume that $A_{2}\neq 0$ is a direct sum of totally positive quaternion algebras. If $A_{1}=0$ then we need to use a slight modification of [BJ11, Algorithm 3.1] as discussed in the further details of Step (11) of Algorithm 8.9.
Let $e_{1},e_{2}\in A$ be the central idempotents such that $A_{i}=e_{i}A$ for $i=1,2$. Let $C$ denote the centre of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order in $C$. Let $\mathfrak{f}$ be any proper full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and let $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$. Then the central idempotents $e_{1}$ and $e_{2}$ induce decompositions
@@
For $i=1,2$ let $T_{i}$ denote the image of $\mathcal{M}_{i}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}_{i}\right)^{\times}$ and note that $T:=T_{1}\times T_{2}$ is the image of ×$\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$.
@@ 12.2. Criteria for certain ideals to be free
    There exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that
@@
The equivalence of (i) and (ii) follows from Proposition 7.1. Assertion (iii) is just an immediate reformulation of (ii). ∎
[BJ11, Algorithm 3.1] essentially uses the the equivalence of conditions (i) and (ii) in Lemma 12.1 to reduce the freeness test to an enumeration over the set $T$ of cardinality $\lvert T\rvert=\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$. In the following, we show how condition (iii) can be used to replace this by an enumeration over $T_{2}$ and computations related to the reduced norm. (See also Remark 12.6.)
@@
${\mathcal{M}_{1}^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\mathrm{nr}}$${\mathcal{O}_{C_{1}}^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\overline{\mathrm{nr}}}$${(\mathcal{O}_{C_{1}}/\mathfrak{g}_{1})^{\times},}$ (12.1)
where the vertical maps are induced by the canonical projections. For $i=1,2$, let $\pi_{i}:\Lambda\longrightarrow\mathcal{M}_{i}$ denote the canonical projection and set $\pi_{i}(\Lambda):=\Lambda_{i}$. Hence we can and do also consider $\pi_{i}$ as a map $\pi_{i}\colon:\Lambda\longrightarrow\Lambda_{i}$, depending on context. Note that in fact $\mathfrak{h}_{i}:=\Lambda\cap\mathcal{M}_{i}$ is a two-sided ideal of $\Lambda_{1}\oplus\Lambda_{2}$ contained in $\Lambda$.
@@
$\overline{\pi}_{i}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda_{i}/\mathfrak{f}_{i})^{\times},$ (12.2)
@@
$S:=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\exists\overline{\gamma}_{2}\in(\Lambda_{2}/\mathfrak{f}_{2})^{\times}:(\overline{\gamma}_{1},\overline{\gamma}_{2})\in\Sigma\right\}.$ (12.3)
@@
$S$ =$=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\overline{\gamma}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}\right\},$
=$=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\gamma_{1},1)\in\Lambda\right\}$
@@
    There exists $\overline{\alpha}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$.
@@
(i) Suppose that $X$ is free over $\Lambda$. Then by Lemma 12.1 there exist $\overline{\alpha}_{2}\in\overline{\beta}_{2}^{-1}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that $(\overline{\alpha}_{1},\overline{\alpha}_{2})\in(\Lambda/\mathfrak{f})^{\times}\text{ and }\overline{\beta}_{1}\overline{\alpha}_{1}\in T_{1}$. Moreover, $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\alpha}_{1})\bmod S\in\mathrm{im}(\sigma)$ by Lemma 12.2, and so we may take $\overline{\gamma}_{1}=\overline{\alpha}_{1}$. The converse holds by Lemmas 12.1 and 12.2.
@@
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\in\mathrm{im}(\sigma)$, return $\mathsf{true}$.
@@
Step 4(b): This can be performed using the commutative diagram (12.1) and standard algorithms for finitely generated abelian groups, as presented for example in [Coh00, §4.1], for example.
@@
Compared to [BJ11, Algorithm 3.1] where the size of the enumeration set is essentially the cardinality of the image of the canonical map ×$\mathcal{M}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$, that is, $\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$, the enumeration set in Algorithm 12.5 is of size $\lvert T_{2}\rvert$, and hence only depends on the component $A_{2}$. Note that $T_{2}$ can be computed as explained in [BJ11, §4.5] (see also [KV10, Remark 7.5]), and this is much faster than computing the image of $\mathcal{M}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ for $\mathcal{M}$ a maximal $\mathcal{O}$-order in an arbitrary finite-dimensional semisimple $K$-algebra. If $|T_{1}|$ is ‘small’, then it may be that [BJ11, Algorithm 3.1] is faster than Algorithm 12.5 thanks to the optimisation described in [BJ11, §7] (also see [HJ20, §9]).
@@ 12.3. Computation of $S$
Recall the definition of $S$ given in (12.3). Since we require only generators of $S$, it will be enough to consider the computation of (generators of) $\Sigma=\ker(\overline{\pi}_{2})$. Recall that $\mathfrak{h}_{1}=\Lambda\cap\mathcal{M}_{1}$ and consider the two-sided ideal $\tilde{\mathfrak{f}}:=\mathfrak{h}_{1}+\mathfrak{f}=\mathfrak{h}_{1}\oplus\mathfrak{f}_{2}$. Let
$\overline{\pi}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda/\tilde{\mathfrak{f}})^{\times}$
@@
    $\Lambda/\mathfrak{f}\cong\prod_{i=1}^{r+s}\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{and}\quad\Lambda/\tilde{\mathfrak{f}}\cong\prod_{i=1}^{r}\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}).$
@@
    where $\tau_{i}\colon(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}))^{\times}$ are the canonical surjections.
@@
$\mathfrak{f}=\bigcap_{i=1}^{r+s}(\mathfrak{q}_{i}\Lambda+\mathfrak{f})\quad\text{ and }\quad\tilde{\mathfrak{f}}=\bigcap_{i=1}^{r}(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}),$
@@
Since we know how to compute generators for each of the groups $\left(\Lambda/(\mathfrak{q}_{r+i}\Lambda+\mathfrak{f})\right)^{\times}$ for $i=1,\ldots,s$ by the results of §11.2, it remains to consider the computation of $\ker(\tau_{i})$ for $i=1,\ldots,r$. Each element in $\ker(\tau_{i})$ is the image in $(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}$ of an element of the form $1+\lambda$ with $\lambda\in(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$. Note, however, that not every element in $1+(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}})$ is actually a unit mod $(\mathfrak{q}_{i}\Lambda+\mathfrak{f})$.
@@
$\tau\colon(\Lambda/(\mathfrak{q}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}\Lambda+\tilde{\mathfrak{f}}))^{\times}$
@@
    $\alpha\text{ is invertible mod }\mathfrak{Q}\iff\alpha\text{ is invertible mod}\mathfrak{Q}_{1}.$
@@
    Identifying $(1+\tilde{\mathfrak{Q}})$ and $(1+\mathfrak{Q}_{1})$ with their images in $\Lambda/\mathfrak{Q}$ we have $\ker(\tau)=(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}$ and the sequence
@@
$1\longrightarrow\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}\longrightarrow\ker(\tau)\longrightarrow\frac{(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}}{1+\mathfrak{Q}_{1}}\longrightarrow 1$ (12.4)
is exact, it suffices to find generators for the first and third group. Generators for $\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}$ can be computed as in [BB06, §3.7]. For the computation of generators of the non-trivial group on the right hand side of (12.4), the elements can be listed as follows: determine a set of representatives of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ and check for each such representative $\alpha$ whether $1+\alpha$ it is a unit mod $\mathfrak{Q}$.

### v3

@@ 11.2. Primary decomposition and generators of unit groups of quotient rings
Let $\mathcal{O}_{C}$ be the (unique) maximal $\mathcal{O}$-order in $C$. This is equal to the integral closure of $\Gamma$ in $C$. By [AM69, Theorem 5.10], each of the maximal ideals $\mathfrak{p}_{i}$ is of the form $\mathfrak{P}_{i}\cap\Gamma$ for some maximal ideal $\mathfrak{P}_{i}$ of $\mathcal{O}_{C}$. Let $\{\mathfrak{P}_{1},\ldots,\mathfrak{P}_{n}\}$ be the set of maximal ideals of $\mathcal{O}_{C}$ with $\mathfrak{g}\mathcal{O}_{C}\subseteq\mathfrak{P}_{i}$. Then
@@
Let $\Lambda$ be an $\mathcal{O}$-order in a finite-dimensional semisimple $K$-algebra AA and let $\mathfrak{h}$ be a full two-sided ideal of $\Lambda$. Then there exists an algorithm that determines generators of $(\Lambda/\mathfrak{h})^{\times}$.
@@
The algorithm is trivial in the case $\mathfrak{h}=\Lambda$, and so henceforth we assume that $\mathfrak{h}\neq\Lambda$. Let $C$ denote the center of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order contained in $C$. Let $\mathfrak{g}=\mathfrak{h}\cap C$. Then $\mathfrak{g}$ is a proper ideal of an $\mathcal{O}$-order $\Gamma$, which may be strictly contained in $\mathcal{O}_{C}$. In the case that $\mathfrak{h}$ is equal to $\mathfrak{f}$, a two-sided ideal of a maximal $\mathcal{O}$-order $\mathcal{M}$ in $A$ containing $\Lambda$, the desired algorithm is described in [BHJ22, §6.6] (this is a minor modification of [BB06, §3.4–3.7]). This relies crucially on the computation of the primary decomposition of $\mathfrak{g}$, which can be accomplished using work of Bley–Endres [BE05] under the assumption that $\Gamma=\mathcal{O}_{C}$ (which must hold when $\mathfrak{h}=\mathfrak{f}$). In fact, the computation of the primary decomposition of $\mathfrak{g}$ is the only obstacle to generalising the algorithm to arbitrary choices of $\mathfrak{h}$. By removing the restriction that $\Gamma=\mathcal{O}_{C}$ on the computation of primary decompositions in Corollary 11.3, we thereby obtain an algorithm for the computation of generators of $(\Lambda/\mathfrak{h})^{\times}$ for arbitrary full two-sided ideals $\mathfrak{h}$ of $\Lambda$. ∎
@@ 12.1. Setup
Henceforth assume that $A$ admits a decomposition $A=A_{1}\oplus A_{2}$ of $K$-algebras such that $A_{1}$ satisfies the Eichler condition. In applications, we can always assume that $A_{2}\neq 0$ is a direct sum of totally definite quaternion algebras. If $A_{1}=0$ then we need to use a slight modification of [BJ11, Algorithm 3.1] as discussed in the further details of Step (11) of Algorithm 8.9.
Let $e_{1},e_{2}\in A$ be the central idempotents such that $A_{i}=e_{i}A$ for $i=1,2$. Let $C$ denote the center of $A$ and let $\mathcal{O}_{C}$ denote the (unique) maximal $\mathcal{O}$-order in $C$. Let $\mathfrak{f}$ be any full two-sided ideal of $\mathcal{M}$ that is contained in $\Lambda$ and let $\mathfrak{g}=\mathfrak{f}\cap\mathcal{O}_{C}$. Then the central idempotents $e_{1}$ and $e_{2}$ induce decompositions
@@
For $i=1,2$ let $T_{i}$ denote the image of $\mathcal{M}_{i}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}_{i}\right)^{\times}$ and note that $T:=T_{1}\times T_{2}$ is the image of $\mathcal{M}^{\times}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$.
@@ 12.2. Criteria for certain ideals to be free
    There exist $\overline{\alpha}_{2}\in\overline{\beta}{}^{-1}_{2}T_{2}$ and $\overline{\alpha}_{1}\in(\mathcal{M}_{1}/\mathfrak{f}_{1})^{\times}$ such that
@@
The equivalence of (i) and (ii) follows from Corollary 7.3. Assertion (iii) is just an immediate reformulation of (ii). ∎
[BJ11, Algorithm 3.1] essentially uses the equivalence of conditions (i) and (ii) in Lemma 12.1 to reduce the freeness test to an enumeration over the set $T$ of cardinality $\lvert T\rvert=\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$. In the following, we show how condition (iii) can be used to replace this by an enumeration over $T_{2}$ and computations related to the reduced norm. (See also Remark 12.6.)
@@
[…]
where the vertical maps are induced by the canonical projections. For $i=1,2$, let $\pi_{i}:\Lambda\longrightarrow\mathcal{M}_{i}$ denote the canonical projection and set $\pi_{i}(\Lambda):=\Lambda_{i}$. Hence we can and do also consider $\pi_{i}$ as a map $\pi_{i}\colon:\Lambda\longrightarrow\Lambda_{i}$, depending on context.
@@
(12.2) $\overline{\pi}_{i}\colon(\Lambda/\mathfrak{f})^{\times}\longrightarrow(\Lambda_{i}/\mathfrak{f}_{i})^{\times},$
@@
(12.3) $S:=\overline{\mathrm{nr}}(\overline{\pi}_{1}(\Sigma))=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\exists\overline{\gamma}_{2}\in(\Lambda_{2}/\mathfrak{f}_{2})^{\times}:(\overline{\gamma}_{1},\overline{\gamma}_{2})\in\Sigma\right\}.$
@@
$S$ $=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\overline{\gamma}_{1},\overline{1})\in\left(\Lambda/\mathfrak{f}\right)^{\times}\right\},$
$=\left\{\overline{\mathrm{nr}}(\overline{\gamma}_{1})\mid\overline{\gamma}_{1}\in\left(\mathcal{M}_{1}/\mathfrak{f}_{1}\right)^{\times}\text{ and }(\gamma_{1},1)\in\Lambda\right\}$
@@
[…]
@@
[…]
@@
      If $\overline{\mathrm{nr}}(\overline{\beta}_{1}\overline{\gamma}_{1})\bmod S\in\mathrm{im}(\sigma)$, return $\mathsf{true}$.
@@
Step 4(b): This can be performed using the commutative diagram (12.1) and standard algorithms for finitely generated abelian groups, as presented in [Coh00, §4.1], for example.
@@
Compared to [BJ11, Algorithm 3.1] where the size of the enumeration set is essentially the cardinality of the image of the canonical map $\mathcal{M}^{\times}\longrightarrow\left(\mathcal{M}/\mathfrak{f}\right)^{\times}$, that is, $\lvert T_{1}\rvert\cdot\lvert T_{2}\rvert$, the enumeration set in Algorithm 12.5 is of size $\lvert T_{2}\rvert$, and hence only depends on the component $A_{2}$. Note that $T_{2}$ can be computed as explained in [BJ11, §4.5] (see also [KV10, Remark 7.5]), and this is much faster than computing the image of $\mathcal{M}\longrightarrow(\mathcal{M}/\mathfrak{f})^{\times}$ for $\mathcal{M}$ a maximal $\mathcal{O}$-order in an arbitrary finite-dimensional semisimple $K$-algebra. If $|T_{1}|$ is ‘small’, then it may be that [BJ11, Algorithm 3.1] is faster than Algorithm 12.5 thanks to the optimisation described in [BJ11, §7] (also see [HJ20, §9]).
@@ 12.3. Computation of $S$
Recall the definition of $S$ given in (12.3). Since we require only generators of $S$, it will be enough to consider the computation of (generators of) $\Sigma=\ker(\overline{\pi}_{2})$. Let $\mathfrak{h}_{1}=\Lambda\cap\mathcal{M}_{1}$ and consider the two-sided ideal $\tilde{\mathfrak{f}}:=\mathfrak{h}_{1}+\mathfrak{f}=\mathfrak{h}_{1}\oplus\mathfrak{f}_{2}$. Let
[…]
@@
[…]
@@
    where $)×\tau_{i}\colon(\Lambda/(\mathfrak{q}_{i}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}_{i}\Lambda+\tilde{\mathfrak{f}}$$)^{\times}$ are the canonical surjections.
@@
[…]
@@
[…]
@@
$)×\tau\colon(\Lambda/(\mathfrak{q}\Lambda+\mathfrak{f}))^{\times}\longrightarrow(\Lambda/(\tilde{\mathfrak{q}}\Lambda+\tilde{\mathfrak{f}}$$)^{\times}$
@@
[…]
@@
[…]
@@
(12.4) $1\longrightarrow\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}\longrightarrow\ker(\tau)\longrightarrow\frac{(1+\tilde{\mathfrak{Q}})\cap\left(\Lambda/\mathfrak{Q}\right)^{\times}}{1+\mathfrak{Q}_{1}}\longrightarrow 1$
is exact, it suffices to find generators for the first and third group. Generators for $\frac{1+\mathfrak{Q}_{1}}{1+\mathfrak{Q}}$ can be computed as in [BB06, §3.7]. For the computation of generators of the nontrivial group on the right hand side of (12.4), the elements can be listed as follows: determine a set of representatives of $\tilde{\mathfrak{Q}}/\mathfrak{Q}_{1}$ and check for each such representative $\alpha$ whether $1+\alpha$ is a unit mod $\mathfrak{Q}$.